# Counting rational points on Hirzebruch–Kleinschmidt varieties over global function fields

### v1

@@ Acknowledgments
The authors thank Giancarlo Lucchini-Arteche for helpful comments on the first Galois cohomology group of an algebraic variety.
@@ 2.1. Basic notation
Let $\mathscr{C}$ be a projective, smooth, geometrically irreducible curve of genus $g$ defined over the finite field $\mathbb{F}_{q}$ of $q$ elements (as usual, $q$ is a positive power of a rational prime). Throughout this article we let $K$ denote the field $\mathbb{F}_{q}(\mathscr{C})$ of rational functions of $\mathscr{C}$. Associated to $K$ we have the following objects (see e.g. [Ros02] for details):
@@
  • For each $v\in\operatorname{Val}(K)$, let $\mathscr{O}_{v}$ be the associated valuation ring and $m_{v}$ its maximal ideal, $k_{v}:=\mathscr{O}_{v}/m_{v}$ and $f_{v}:=[k_{v}:\mathbb{F}_{q}]$, so we have #$\#k_{v}=q^{f_{v}}$. We define for $x\in K$, $|x|_{v}:=q^{-f_{v}v(x)}$, so that for every $x\in K,x\neq 0$ we have the product formula
@@ 2.2. Geometric and arithmetic tools
As a consequence of the Riemann–Roch theorem for curves (see e.g. [Ros02, Theorem 5.4]) we have that
@@
$\operatorname{Z}_{K}(T):=\sum_{D\geq 0}T^{\deg(D)}=\prod_{v\in\operatorname{Val}(K)}\left(1-T^{f_{v}}\right),$
@@
The series defining $\zeta_{K}(s)$ converges absolutely and uniformly on compact subsets of the half-plane $\Re(s)>1$ and has no zeroes in that domain. Moreover, it is a rational function in $q^{-s}$ and has a simple pole at $s=1$ with residue
@@
(see e.g. [Ros02, Theorem 5.9]).
@@ 3. Hirzebruch–Kleinschmidt varieties
Let $X$ be a smooth projective toric variety with $\operatorname{Pic}(X)\simeq\mathbb{Z}^{2}$. Then, there exists integers $r\geq 1,t\geq 2,0\leq a_{1}\leq\cdots\leq a_{r}$ with $r+t-1=d=\dim(X)$ such that
@@ 3.1. Effective divisors
In this section we give a basis of $\operatorname{Pic}(X)$ for a Hirzebruch–Kleinschmidt variety $X$, and use it to describe the anticanonical class, the effective cone, and the big divisor classes.
@@
was given in [HMM24, Proposition 3.4]. More precisely, putting $f_{Y}:=\pi_{Y}^{\ast}\mathscr{O}_{\mathbb{P}^{t-1}}(1)$ and $h_{Y}:=\mathscr{O}_{Y}(1)$, where $\pi_{Y}:Y\to\mathbb{P}^{t-1}$ is the natural projection, it was proven that
@@
the line bundle classes $-K_{Y},f_{Y}$ and $h_{Y}$ correspond to $-K_{X},f$ and $h-a_{r}f$, respectively (see e.g. [Har77, Chapter V, Corollary 2.18]). Then, simple computations lead to the desired results. We omit the details for brevity. ∎
It follows from [Laz04, Theorem 2.2.26] that a divisor class in $\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$ is big if and only if it lies in the interior of the effective cone $\Lambda_{\textup{eff}}(X)$. Hence, we get the following corollary from Proposition 3.2.
@@ 4. Height zeta function of the projective space
$H_{\mathbb{P}^{n}}([x_{0},\ldots,x_{n}]):=\prod_{v\in\operatorname{Val}(K)}\sup\{|x_{0}|_{v},|x_{1}|_{v},\ldots,|x_{n}|_{v}\},$ (9)
@@
$\operatorname{Res}_{s=n+1}\zeta_{\mathbb{P}^{n}}(s)=\frac{h_{K}q^{(n+1)(1-g)}}{\zeta_{K}(n+1)(q-1)\log(q)}.$
Proof.
The fact that $\zeta_{\mathbb{P}^{n}}(s)$ is a rational function in $q^{-s}$ follows from [Wan92, Theorem 3.2]. Now, for $m\geq 1$ put $N(m):=\#\{P\in\mathbb{P}^{n}(K):H(P)=m\}$. By definition of the standard height function, we have $N(m)=0$ if $m$ is not of the form $m=q^{d}$ with $d\geq 0$ an integer. Hence
$\zeta_{\mathbb{P}^{n}}(s)=\sum_{d\geq 0}N(q^{d})q^{-ds}.$
Now, by [Ser89, Section 2.5] we have the estimate
$N(q^{d})=\frac{h_{K}q^{(n+1)(1-g)}}{\zeta_{K}(n+1)(q-1)}q^{d(n+1)}+O(q^{dn})\quad\text{as }d\to\infty.$ (10)
This implies that $\zeta_{\mathbb{P}^{n}}(s)$ converges absolutely for $\Re(s)>n+1$, and moreover
@@
with $F(s)$ an analytic function on $\Re(s)>n$. Hence, we also get
@@
Remark 4.2.
  1. (1)
    A better estimate for the error term in (10) is given in [Wan92, Corollary 4.3], which is related to the fact that $\zeta_{\mathbb{P}^{n}}(s)$ is analytic for $\Re(s)\in\left]\frac{1}{2},n+1\right[$. For the purpose of this paper, however, we do not require such a strong result.
  2. (2)
    Theorem 4.1 also follows from [Pey12, Théorème 3.11].
@@ 5. Height zeta functions of Hirzebruch–Kleinschmidt varieties
In this section, we use the explicit equations defining our Hirzebruch–Kleinschmidt varieties in order to study the analytic properties of their height zeta functions. To do this, we first present some lemmas concerning Möbius inversion, supremum of divisors and relations between different zeta functions. We will first focus on the case $a_{r}>0$ in the family of varieties $X=X_{d}(a_{1},\ldots,a_{r})$, and later treat the much easier case where $a_{r}=0$, since then $X\simeq\mathbb{P}^{r}\times\mathbb{P}^{t-1}$ and the relevant height zeta functions are just products of height zeta functions of projective spaces.
@@
As explained in the introduction, to carry out the analysis of the corresponding height zeta functions, we find it necessary to remove a closed subset in order to avoid accumulation of rational points of bounded height. To this purpose, the following definition was given in the introduction.
@@ 5.1. Height zeta functions
$A_{L}:=\frac{r+1}{\gamma},\quad B_{L}:=\frac{(r+1)a_{r}-|\mathbf{a}|+t}{\gamma a_{r}+\xi}.$ (12)
@@
We now present the first main result of this section, which describes the analytical properties of the function $\zeta_{U,L}(s)$ in the case where $a_{r}>0$.
@@
Let $X:=X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety over the global function field $K=\mathbb{F}_{q}(\mathscr{C})$ with $a_{r}>0$. Moreover, let $U:=U_{d}(a_{1},\ldots,a_{r})$ be the good open subset of $X$, let $L=\gamma h+\xi f$ be a big line bundle class in $\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$, and let $\zeta_{U,L}(s)$ be the associated height zeta function. Then, $\zeta_{U,L}(s)$ is a rational function in $q^{-s}$. Moreover, $\zeta_{U,L}(s)$ converges absolutely for $\Re(s)>a(L)$ and it has a pole of order $b(L)$ at $s=a(L)$ with
$\lim_{s\to a(L)}(s-a(L))^{b(L)}\zeta_{U,L}(s)=\left\{\begin{array}[]{ll}\vspace{0.2cm}\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)(\gamma a_{r}+\xi)\gamma(q-1)^{2}\log(q)^{2}}&\text{if }\xi=\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma,\\ \vspace{0.2cm}\frac{q^{(d+2-N_{X})(1-g)}h_{K}\mathbf{R}_{K}(1-N_{X},\gamma B_{L}-r+N_{X}-1)}{\zeta_{K}\left(t\right)\zeta_{K}(\gamma B_{L})(\gamma a_{r}+\xi)(q-1)\log(q)}&\text{if }\xi<\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma,\\ \vspace{0.2cm}\frac{q^{(r+1)(1-g)}h_{K}\mathbf{R}_{K}\left(1-t,A_{L}\xi+|\mathbf{a}|\right)}{\zeta_{K}\left(A_{L}\xi+|\mathbf{a}|\right)\zeta_{K}(r+1)\gamma(q-1)\log(q)}&\text{if }\xi>\left(\frac{t-|\mathbf{a}|}{r+1}\right)\gamma,\\ \end{array}\right.$
where $N_{X}:=\#\{j\in\{1,\ldots,r\}:a_{j}=a_{r}\}$ and

### v2

@@ Acknowledgments
The authors thank Loïs Faisant and Daniel Loughran for geometric insight on the Manin–Peyre conjecture for varieties defined over global function fields, and Giancarlo Lucchini-Arteche for helpful comments on the first Galois cohomology group of an algebraic variety.
@@ 2.1. Basic notation
Let $\mathscr{C}$ be a projective, smooth, geometrically irreducible curve of genus $g$ defined over the finite field $\mathbb{F}_{q}$ of $q$ elements (as usual, $q$ is a positive power of a rational prime). Throughout this article we let $K$ denote the field $\mathbb{F}_{q}(\mathscr{C})$ of rational functions of $\mathscr{C}$. Associated to $K$ we have the following objects (see, e.g., [Ros02] for details):
@@
  • For each $v\in\operatorname{Val}(K)$, let $\mathscr{O}_{v}$ be the associated valuation ring and $m_{v}$ its maximal ideal, $k_{v}:=\mathscr{O}_{v}/m_{v}$ and $f_{v}:=[k_{v}:\mathbb{F}_{q}]$, so we have #$\#k_{v}=q^{f_{v}}$. We define for $x\in K$, $|x|_{v}:=q^{-f_{v}v(x)}$, so that for every $x\in K^{\times}$ we have the product formula
@@ 2.2. Geometric and arithmetic tools
As a consequence of the Riemann–Roch theorem for curves (see, e.g., [Ros02, Theorem 5.4]) we have that
@@
$\operatorname{Z}_{K}(T):=\sum_{D\geq 0}T^{\deg(D)}=\prod_{v\in\operatorname{Val}(K)}\left(1-T^{f_{v}}\right)^{-1},$
@@
The series defining $\zeta_{K}(s)$ converges absolutely and uniformly on compact subsets of the half-plane $\Re(s)>1$ and has no zeroes in that domain. Moreover, there exits a polynomial $\mathbf{L}_{K}(X)\in\mathbb{Z}[X]$ of degree $2g$ such that
$\zeta_{K}(s)=\frac{\mathbf{L}_{K}(q^{-s})}{(1-q^{-s})(1-q^{1-s})}$
and $\zeta_{K}(s)$ has simple poles at $s=1$ and $s=0$ with
@@
(see, e.g., [Ros02, Theorem 5.9]).
2.3. A Tauberian theorem for global function fields
The following theorem is a variant of [Ros02, Theorem 17.4].
Theorem 2.1.
Let $X$ be a countable set and $H:X\to q^{\mathbb{N}_{0}}$ a function. Suppose that
$\zeta_{H}(s):=\sum_{x\in X}H(x)^{-s}$
is absolutely convergent for $\Re(s)>a>0$ and there exists $\varepsilon>0$ such that $\zeta_{H}(s)$ extends to a meromorphic function on the half-plane $\Re(s)\geq a-\varepsilon$ satisfying the following properties:
  1. (1)
    $\zeta_{H}(s)$ has a pole of order $b\geq 1$ at $s=s_{0}:=a$,
  2. (2)
    There exists real numbers $10=\theta_{0}<\theta_{1}<\ldots<\theta_{n}<$ 1 such that $\zeta_{H}(s)$ is holomorphic for $\Re(s)>a-\varepsilon,s\not\in\left\{a+\frac{2\pi i\theta_{k}}{\log(q)}+\frac{2\pi im}{\log(q)}:k\in\{0,\ldots,n\},m\in\mathbb{Z}\right\}$.
  3. (3)
    For every $k\in\{1,\ldots,n\}$, $\zeta_{H}(s)$ is either holomorphic or has a pole of order $\leq b$ at $s=s_{k}:=a+\frac{2\pi i\theta_{k}}{\log(q)}$.
Then, for every $M\geq 0$ the cardinality
$\tilde{N}(X,H,q^{M}):=\#\{x\in X:H(x)=q^{M}\}$
is finite and for each $k\in\{0,\ldots,n\}$ there exists a polynomial $Q_{k}(M)$ of degree $b-1$ and leading coefficient $r_{k}:=\lim_{s\to s_{k}}(s-s_{k})^{b}\zeta_{H}(s)$ such that
$\tilde{N}(X,H,q^{M})=\frac{\log(q)^{b}}{(b-1)!}q^{aM}\sum_{k=0}^{n}e^{2\pi i\theta_{k}M}Q_{k}(M)+O\left(q^{\delta M}\right)\quad\text{as }M\to\infty,M\in\mathbb{Z},$
for every $\delta>a-\varepsilon$.
Proof.
For $\Re(s)>a$ we have
$\zeta_{H}(s)=\sum_{M=0}^{\infty}\tilde{N}(X,H,q^{M})q^{-sM}.$
This implies that $\tilde{N}(X,H,q^{M})$ is finite for every $M\geq 0$. Now, put
$\operatorname{Z}_{H}(T):=\sum_{M=0}^{\infty}\tilde{N}(X,H,q^{M})T^{M}$ (8)
so that $\operatorname{Z}_{H}(q^{-s})=\zeta_{H}(s)$. By hypothesis, the series $\operatorname{Z}_{H}(T)$ converges absolutely for $|T|<q^{-a}$ and it extends to a meromorphic function on $|T|\leq q^{-(a-\varepsilon)}$ with poles contained in $|T|\in\{q^{-a},q^{-(a-\varepsilon)}\}$. Since the closed disk $|T|\leq q^{-(a-\varepsilon)}$ is compact, there exists $\varepsilon^{\prime}>\varepsilon$ such that $\operatorname{Z}_{H}(T)$ is meromorphic for $|T|\leq q^{-(a-\varepsilon^{\prime})}$ and has no poles on $q^{-(a-\varepsilon)}<|T|\leq q^{-(a-\varepsilon^{\prime})}$. Fix $\eta\in\,]0,q^{-a}[$ and consider the curves $\gamma_{1}(t):=\eta e^{it}$ and $\gamma_{2}(t):=q^{-(a-\varepsilon^{\prime})}e^{it}$ for $t\in[0,2\pi]$. Recall that $s_{k}:=a+\frac{2\pi i\theta_{k}}{\log(q)}$ and put $T_{k}:=q^{-s_{k}}$ for $k\in\{0,\ldots,n\}$. Then, $T_{0},\ldots,T_{n}$ are exactly the poles of $\operatorname{Z}_{H}(T)$ in $|T|=q^{-a}$. We also denote by $\{T_{n+1},\ldots,T_{n+m}\}$ the (possibly empty) set of poles of $\operatorname{Z}_{H}(T)$ in $|T|=q^{-(a-\varepsilon)}$. Now, for every integer $M>0$ we have, by Cauchy’s residue theorem, the equality
$\frac{1}{2\pi i}\int_{\gamma_{2}}\frac{\operatorname{Z}_{H}(T)}{T^{M+1}}\textup{d}t=\frac{1}{2\pi i}\int_{\gamma_{1}}\frac{\operatorname{Z}_{H}(T)}{T^{M+1}}\textup{d}t+\sum_{k=0}^{n+m}\operatorname{Res}_{T=T_{k}}\left(\frac{\operatorname{Z}_{H}(T)}{T^{M+1}}\right)$ (9)
On the one hand, since $\operatorname{Z}_{H}(T)$ is holomorphic for $|T|=q^{-(a-\varepsilon^{\prime})}$, we have
$\frac{1}{2\pi i}\int_{\gamma_{2}}\frac{\operatorname{Z}_{H}(T)}{T^{M+1}}\textup{d}t=O\left(q^{(a-\varepsilon^{\prime})M}\right)$ (10)
with an implicit constant independent of $M$. On the other hand, from (8) we have
$\frac{1}{2\pi i}\int_{\gamma_{1}}\frac{\operatorname{Z}_{H}(T)}{T^{M+1}}\textup{d}t=\tilde{N}(X,H,q^{M}).$ (11)
Next, for $k\in\{0,\ldots,n\}$ we write
$\operatorname{Z}_{H}(T)=\sum_{j=-b}^{\infty}c_{k,j}(T-T_{k})^{j}$
with $c_{k,j}\in\mathbb{C}$ and recall the Taylor expansion
$T^{-(M+1)}=\sum_{j=0}^{\infty}{M+j\choose M}(-1)^{j}T_{k}^{-(M+j+1)}(T-T_{k})^{j},$
in order to get
$\operatorname{Res}_{T=T_{k}}\left(\frac{\operatorname{Z}_{H}(T)}{T^{M+1}}\right)$ $=$ $-T_{k}^{-M}\sum_{j=1}^{b}c_{k,-j}{M+j-1\choose M}(-1)^{j}T_{k}^{-j}$
$=$ $-q^{Ms_{k}}\sum_{j=1}^{b}c_{k,-j}{M+j-1\choose M}(-1)^{j}q^{js_{k}}.$
Since (${M+j-1\choose M}$ is a polynomial in $M$ of degree $j-1$ and leading term $\frac{M^{j-1}}{(j-1)!}$, we obtain
$\operatorname{Res}_{T=T_{k}}\left(\frac{\operatorname{Z}_{H}(T)}{T^{M+1}}\right)=(-1)^{b+1}q^{Ms_{k}}\frac{P_{k}(M)}{(b-1)!}q^{bs_{k}},$
where $P_{k}(M)$ is a polynomial of degree $b-1$ with leading coefficient $c_{k,-b}$. But
$c_{k,-b}$ $=$ $\lim_{T\to T_{k}}(T-T_{k})^{b}\operatorname{Z}_{H}(T)$
$=$ $\lim_{s\to s_{k}}(q^{-s}-q^{-s_{k}})^{b}\zeta_{H}(s)$
$=$ $\lim_{s\to s_{k}}\left(\frac{q^{-s}-q^{-s_{k}}}{s-s_{k}}\right)^{b}\lim_{s\to s_{k}}(s-s_{k})^{b}\zeta_{H}(s)$
$=$ $(-\log(q)q^{-s_{k}})^{b}r_{k},$
hence
$\operatorname{Res}_{T=T_{k}}\left(\frac{\operatorname{Z}_{H}(T)}{T^{M+1}}\right)=-q^{Ms_{k}}\log(q)^{b}\frac{Q_{k}(M)}{(b-1)!}=-q^{Ma}e^{2\pi i\theta_{k}M}\log(q)^{b}\frac{Q_{k}(M)}{(b-1)!},$ (12)
where $Q_{k}(M)$ is a polynomial of degree $b-1$ with leading coefficient $r_{k}$. Similarly, for $k\in\{n+1,\ldots,n+m\}$ we have
$\operatorname{Res}_{T=T_{k}}\left(\frac{\operatorname{Z}_{H}(T)}{T^{M+1}}\right)=O\left(M^{c-1}q^{M(a-\varepsilon)}\right),$ (13)
for every integer $c\geq 1$ larger than the maximum among the order of the poles of $\operatorname{Z}_{H}(s)$ on $|T|=q^{-(a-\varepsilon)}$ (if $\operatorname{Z}_{H}(s)$ has no poles in that circle, then we choose $c=1$). Then, the desired results follows from (9), (10), (11), (12) and (13), together with the fact that
$M^{c-1}q^{M(a-\varepsilon)}=O\left(q^{\delta M}\right)\quad\text{as }M\to\infty,$
for every $\delta>a-\varepsilon$. This proves the theorem. ∎
Remark 2.2.
  1. (1)
    The third condition in Theorem 2.1 is equivalent to the existence of the limit $r_{k}=\lim_{s\to s_{k}}(s-s_{k})^{b}\zeta_{H}(s)$ for every $k\in\{1,\ldots,n\}$. In particular, there is no need to compute the exact order of $\zeta_{H}(s)$ at each point $s_{k}$.
  2. (2)
    In all our applications of Theorem 2.1, the points $s_{k}=a+\frac{2\pi i\theta_{k}}{\log(q)}$ can be assumed to have $\theta_{k}=\frac{k}{N}$ with $k=0,\ldots,N-1$ for some integer $N\geq 1$. This implies
    $\tilde{N}(X,H,q^{M})=\frac{\log(q)^{b}}{(b-1)!}Q(M)q^{aM}+O\left(q^{\delta M}\right)\quad\text{as }M\to\infty,M\in\mathbb{Z},$
    with
    $Q(M):=C_{b-1}(M)M^{b-1}+C_{b-2}(M)M^{b-2}+\ldots+C_{0}(M)$
    and $C_{k}(M)\in\mathbb{R}$ depending only on the residue class of $M$ mod $N$. Moreover, the leading term has
    $C_{b-1}(M)=\sum_{k=0}^{N-1}e^{2\pi ikM/N}r_{k}=N\lim_{s\to a^{+}}(s-a)^{b}\sum_{\begin{subarray}{c}m\geq 0\\ m\equiv M\,(\text{mod }N)\end{subarray}}\tilde{N}(X,H,q^{m})q^{-sm}.$
    In particular, if $C_{b-1}(M_{0})>0$ for a given integer $M_{0}\in\{0,\ldots,N-1\}$, then one obtains the asymptotic formula
    $\tilde{N}(X,H,q^{M})\sim\left(\frac{\log(q)^{b}}{(b-1)!}C_{b-1}(M_{0})\right)M^{b-1}q^{aM}\quad\text{as }M\to\infty,M\equiv M_{0}\,(\text{mod }N).$
    Finally, note that the first condition in the above theorem implies that $C_{b-1}(M_{0})>0$ for at least one $M_{0}\in\{0,\ldots,N-1\}$.
@@ 3. Hirzebruch–Kleinschmidt varieties
Let $X$ be a smooth projective toric variety with $\operatorname{Pic}(X)\simeq\mathbb{Z}^{2}$. Then, there exist integers $r\geq 1,t\geq 2,0\leq a_{1}\leq\cdots\leq a_{r}$ with $r+t-1=d=\dim(X)$ such that
@@ 3.1. Effective divisors
In this section we give a basis of $\operatorname{Pic}(X)$ for a Hirzebruch–Kleinschmidt variety $X$, and use it to describe the anticanonical class, the effective cone, and the big line bundle classes.
@@
is given in [HMM24, Proposition 3.4]. More precisely, putting $f_{Y}:=\pi_{Y}^{\ast}\mathscr{O}_{\mathbb{P}^{t-1}}(1)$ and $h_{Y}:=\mathscr{O}_{Y}(1)$, where $\pi_{Y}:Y\to\mathbb{P}^{t-1}$ is the natural projection, it is proven that
@@
the line bundle classes $-K_{Y},f_{Y}$ and $h_{Y}$ correspond to $-K_{X},f$ and $h-a_{r}f$, respectively (see, e.g., [Har77, Chapter V, Corollary 2.18]). Then, simple computations lead to the desired results. We omit the details for brevity. ∎
It follows from [Laz04, Theorem 2.2.26] that a line bundle class in $\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$ is big if and only if it lies in the interior of the effective cone $\Lambda_{\textup{eff}}(X)$. Hence, we get the following corollary from Proposition 3.2.
@@ 4. Height zeta function of the projective space
$H_{\mathbb{P}^{n}}([x_{0}:\ldots:x_{n}]):=\prod_{v\in\operatorname{Val}(K)}\sup\{|x_{0}|_{v},|x_{1}|_{v},\ldots,|x_{n}|_{v}\},$ (15)
@@
$\operatorname{Res}_{s=n+1}\zeta_{\mathbb{P}^{n}}(s)=\frac{h_{K}q^{(n+1)(1-g)}}{\zeta_{K}(n+1)(q-1)\log(q)},$
and it is holomorphic in $\Re(s)>\frac{1}{2},s\not\in\left\{n+1+\frac{2\pi im}{\log(q)}:m\in\mathbb{Z}\right\}$.
In the proof of Theorem 4.1 below, we use the following estimate for
$\tilde{N}(\mathbb{P}^{n},q^{d}):=\#\{P\in\mathbb{P}^{n}(K):H_{\mathbb{P}^{n}}(P)=q^{d}\}$
due to Wan ([Wan92, Corollary 4.3] [^5] There is a misprint in the exponent of the error term in [Wan92, Equation (4.8)]. By [Wan92, Theorem 4.1] this error term is bounded above by $q^{\frac{d}{2}}$ times a polynomial in $d$.): For every $\varepsilon>0$ we have
$\tilde{N}(\mathbb{P}^{n},q^{d})=\frac{h_{K}q^{(n+1)(1-g)}}{\zeta_{K}(n+1)(q-1)}q^{d(n+1)}+O\left(q^{d\left(\frac{1}{2}+\varepsilon\right)}\right)\quad\text{as }d\to\infty,d\in\mathbb{Z}.$ (16)
Proof of Theorem 4.1.
The fact that $\zeta_{\mathbb{P}^{n}}(s)$ is a rational function in $q^{-s}$ follows from [Wan92, Theorem 3.2]. Now, by definition of the standard height function (15), we have
$\zeta_{\mathbb{P}^{n}}(s)=\sum_{d\geq 0}\tilde{N}(\mathbb{P}^{n},q^{d})q^{-ds}.$
Then, (16) implies that $\zeta_{\mathbb{P}^{n}}(s)$ converges absolutely for $\Re(s)>n+1$, and moreover
@@
with $F(s)$ an analytic function for $\Re(s)>\frac{1}{2}$. In particular, $\zeta_{\mathbb{P}^{n}}(s)$ is analytic in $\Re(s)>\frac{1}{2},s\not\in\left\{n+1+\frac{2\pi im}{\log(q)}:m\in\mathbb{Z}\right\}$. We also get
@@ 5. Height zeta functions of Hirzebruch–Kleinschmidt varieties
In this section, we use the explicit equations (14) defining our Hirzebruch–Kleinschmidt varieties in order to study the analytic properties of their height zeta functions. To do this, we first present some lemmas concerning Möbius inversion, supremum of divisors and relations between different zeta functions. We will first focus on the case $a_{r}>0$ in the family of varieties $X=X_{d}(a_{1},\ldots,a_{r})$, and later treat the much easier case where $a_{r}=0$, since then $X\simeq\mathbb{P}^{r}\times\mathbb{P}^{t-1}$ and the relevant height zeta functions are just products of height zeta functions of projective spaces.
@@
As explained in the introduction, to carry out the analysis of the corresponding height zeta functions, we find it necessary to remove a closed subset in order to avoid convergence issues, due to accumulation of rational points of bounded height. For this purpose, the following definition was given in the introduction.
@@ 5.1. Height zeta functions
$A_{L}:=\frac{r+1}{\gamma},\quad B_{L}:=\frac{(r+1)a_{r}-|\mathbf{a}|+t}{\gamma a_{r}+\xi},$ (18)
and put
$\mathscr{A}_{L}$ ::=:= $\left\{A_{L}+\frac{2\pi ik}{\gamma\log(q)}:k\in\{1,\ldots,\gamma-1\}\right\},$
$\mathscr{B}_{L}$ ::=:= $\left\{B_{L}+\frac{2\pi ik}{(\gamma a_{r}+\xi)\log(q)}:k\in\{1,\ldots,\gamma a_{r}+\xi-1\}\right\}.$
@@
Below we present the first main result of this section, which describes the analytic properties of the function $\zeta_{U,L}(s)$ in the case where $a_{r}>0$. First, we note that if $L\in\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$ is big and $m\geq 1$ is an integer, then $mL$ is also big and $H_{mL}=H_{L}^{m}$. Moreover, we have the equality of height zeta functions $\zeta_{U,mL}(s)=\zeta_{U,L}(ms)$. In particular, the analytic properties of $\zeta_{U,mL}(s)$ can be deduced from those of $\zeta_{U,L}(s)$. This shows that we can restrict ourselves to the case where $L$ is primitive in $\operatorname{Pic}(X_{d}(a_{1},\ldots,a_{r}))$. Writing $L=\gamma h+\xi f$, this is equivalent to the condition $\mathrm{g.c.d.}(\gamma,\xi)=1$.
In order to state the next theorem, we define
Note that the Riemann–Roch theorem (see Section 2.2) implies the equality
$\mathbf{R}_{K}(a,b)=\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}q^{-(a\ell(D)+b\deg(D))}+q^{a(g-1)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)>2g-2\end{subarray}}q^{-(a+b)\deg(D)}.$ (20)
Hence, $\mathbf{R}_{K}(a,b)=q^{a(g-1)}\zeta_{K}(a+b)+\mathbf{S}_{K}(a,b)$ with the finite sum
In particular, $\mathbf{R}_{K}(a,b)=q^{-a}\zeta_{K}(a+b)$ if $g=0$.
@@
Let $X:=X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety over the global function field $K=\mathbb{F}_{q}(\mathscr{C})$ with $a_{r}>0$. Moreover, let $U:=U_{d}(a_{1},\ldots,a_{r})$ be the good open subset of $X$, let $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ be a big and primitive line bundle class and let $\zeta_{U,L}(s)$ be the associated height zeta function. Then, $\zeta_{U,L}(s)$ converges absolutely for $\Re(s)>a(L)$ and is a rational function in $q^{-s}$. Moreover, the following properties hold:
  1. (1)
    If $A_{L}=B_{L}$, then $\zeta_{U,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{A_{L}-\frac{1}{\gamma},B_{L}-\frac{1}{\gamma a_{r}+\xi}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{A}_{L}\cup\mathscr{B}_{L},m\in\mathbb{Z}\right\},$
    it has a pole of order $b(L)=2$ at $s=a(L)$ and satisfies
    $\lim_{s\to w}(s-w)^{2}\zeta_{U,L}(s)=\left\{\begin{array}[]{ll}\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)(\gamma a_{r}+\xi)\gamma(q-1)^{2}\log(q)^{2}}&\text{if }w=a(L),\\ 0&\text{if }w\in\mathscr{A}_{L}\cup\mathscr{B}_{L}.\end{array}\right.$
  2. (2)
    If $A_{L}<B_{L}$, then $\zeta_{U,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{A_{L},B_{L}-\frac{1}{\gamma a_{r}+\xi}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{B}_{L},m\in\mathbb{Z}\right\},$
    it has a pole of order $b(L)=1$ at $s=a(L)$ and for every $w\in\{a(L)\}\cup\mathscr{B}_{L}$ it satisfies
    $\lim_{s\to w}(s-w)\zeta_{X,L}(s)=\frac{q^{(d+2-N_{X})(1-g)}h_{K}\mathbf{R}_{K}(1-N_{X},\gamma w-r+N_{X}-1)}{\zeta_{K}\left(t\right)\zeta_{K}(\gamma w)(\gamma a_{r}+\xi)(q-1)\log(q)},$
    where $N_{X}:=\#\{j\in\{1,\ldots,r\}:a_{j}=a_{r}\}$.
  3. (3)
    If $A_{L}>B_{L}$, then $\zeta_{U,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{A_{L}-\frac{1}{\gamma},B_{L}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{A}_{L},m\in\mathbb{Z}\right\},$
    it has a pole of order $b(L)=1$ at $s=a(L)$ and for every $w\in\{a(L)\}\cup\mathscr{A}_{L}$ it satisfies
    $\lim_{s\to w}(s-w)\zeta_{U,L}(s)=\frac{q^{(r+1)(1-g)}h_{K}\mathbf{R}_{K}\left(1-t,(\gamma a_{r}+\xi)w-((r+1)a_{r}-|\mathbf{a}|)\right)}{\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)w-\left((r+1)a_{r}-|\mathbf{a}|\right)\right)\zeta_{K}(r+1)\gamma(q-1)\log(q)}.$